# Low-Luminosity Type IIP Supernovae from the Zwicky Transient Facility Census of the Local Universe. II: Lightcurve Analysis

### v1

@@ 1 Introduction
Low-luminosity Type IIP supernovae (LLIIP SNe), defined here as those with mag, as motivated by prior literature (e.g., Spiro et al., 2014), are particularly intriguing because they are thought to originate from progenitor stars with just enough mass to undergo core collapse. These events likely arise from 8–12 progenitors, as supported by pre-SN progenitor imaging (Maund et al., 2005; Li et al., 2006; Mattila et al., 2008; O’Neill et al., 2019; Van Dyk et al., 2023), light curve simulations (Pumo et al., 2017; Fraser et al., 2011), evolutionary models of low-mass red supergiants (RSGs) (Lisakov et al., 2017), and nebular spectroscopy (Jerkstrand et al., 2018). They provide valuable insight into the lower mass threshold for core collapse, which defines the boundary between stars that form white dwarfs and those that form neutron stars or black holes. LLIIP SNe also serve as key candidates for probing the upper mass limit of super-asymptotic giant branch (sAGB) star formation, exploring whether they can explode as electron-capture supernovae (ECSN) (Nomoto, 1984), and understanding their role in shaping elemental abundances in subsequent generations of stars (Tsujimoto & Shigeyama, 1999; Siess & Pumo, 2006). Furthermore, a typical initial mass function (IMF) predicts that more than 40% of CCSN progenitors fall within the 8–12 range, with 25% within the 8–10 subset (Sukhbold et al., 2016). It has also been suggested that LLIIP SNe may instead originate from more massive RSGs with significant fallback (e.g., Zampieri et al., 2003). Thus, systematic studies of LLIIP SNe are essential for improving our understanding of the progenitors, explosion properties, and the dust and chemical enrichment contributed by nearly half of all massive stars that undergo core-collapse.
However, to date, no study of the lightcurve properties of LLIIP SNe has been conducted using data from a systematic SN survey. Due to their low luminosity, they are relatively difficult to detect and classify. There are about a dozen such objects presented in the literature (Turatto et al., 1998; Spiro et al., 2014; Jäger et al., 2020; Müller-Bravo et al., 2020; Reguitti et al., 2021; Yang et al., 2021; Valerin et al., 2022; Kozyreva et al., 2022; Bostroem et al., 2023; Teja et al., 2024; Dastidar et al., 2025). In Part I of the series of three papers on LLIIP SNe (Das et al., 2025), we measured the luminosity function and volumetric rate of 36 LLIIP SNe (and 320 Type IIP SNe) from the Zwicky Transient Facility Census of the Local Universe (ZTF CLU; De et al., 2020) survey. This paper delves into their lightcurve properties.
@@
One of the most used methods to analyze the progenitor and explosion properties is modeling of SN lightcurves. It is well known that lightcurves are sensitive to the physical properties of their progenitors (final masses and radii), as well as the properties of the explosion itself (released energy, amount of synthesized radioactive nickel and its distribution). In this paper, we analyze the observed properties of the multi-band lightcurves of 16 LLIIP SNe. We also infer physical explosion parameters using semi-analytical and radiation-hydrodynamical models that link SN observables to progenitor properties (e.g., Litvinova & Nadezhin, 1983; Hamuy, 2003; Kasen & Woosley, 2009; Goldberg et al., 2019; Morozova et al., 2018; Moriya et al., 2023). Such models have been previously applied to both individual events and larger samples of Type II SNe (Hamuy, 2003; Arcavi et al., 2011; Anderson et al., 2014; Spiro et al., 2014; Sanders et al., 2015; Valenti et al., 2016; Gutiérrez et al., 2017; Morozova et al., 2018; Martinez et al., 2022a, b; Subrayan et al., 2023; Silva-Farfán et al., 2024; Fang et al., 2025). To place these results in a broader context, we also present the lightcurve analysis of the entire sample of 129 Type IIP SNe, obtained from the largest systematic volume-limited SN survey to date.
This paper is structured as follows: in Section 2, we define the sample selection criteria. The data and extinction correction methods are described in Sections 3 and 4, respectively. Sections 5 and 6 detail the fitting methods and results. We discuss the implications of the derived physical properties in Section 7 and conclude in Section 8.
2 Sample Selection
@@
We refer to Paper I (Das et al., 2025) for the complete sample selection criteria to select 330 Type IIP SNe from the ZTF (Bellm et al., 2019; Graham et al., 2019; Dekany et al., 2020; Masci et al., 2019) CLU experiment. To select the sample for lightcurve analysis, we require that the end of the plateau phase is observed in the -band lightcurve. This is to ensure we can extract reliable progenitor and explosion parameters from the lightcurves. The epoch of the end of the plateau () is determined as the epoch where the derivative of the slope of the -band lightcurve () reach a local minimum, and its value is mag day-2, indicating a sharp transition (see Section 5.1 and Figure 1 for details). This marks the end of the plateau phase in the lightcurve evolution. Using this criterion, we identified 129 candidates that fulfill this selection criterion. Out of these, 68 objects have -band data points in the radioactive tail phase, while 61 SNe do not have an observed radioactive tail.
3 Data
@@
3.1 Optical photometry
We perform forced photometry on the ZTF difference images using the ZTF forced-photometry service developed by Masci et al. (2019) in , and bands. For this work, we consider anything less than a 3 detection an upper limit. In the tail phase, for a few SNe we also took photometry in the , , bands with the Optical Imager (IO:O) at the 2.0 m robotic Liverpool Telescope (Steele et al., 2004) and the Alhambra Faint Object Spectrograph and Camera at the Nordic Optical Telescope (Djupvik & Andersen, 2010). The P60 and LT data sets were processed with the FPipe image subtraction pipeline (Fremling et al., 2016), using Sloan Digital Sky Survey (SDSS; Ahn et al., 2012) and PanSTARRS (PS1; Chambers et al., 2016) reference frames. See Das et al. (2025) for details on photometric observations and reductions. Figure 1 shows a few representative lightcurves as examples. The photometry data and lightcurve plots for all the SNe in the sample will be made available on Zenodo.
3.2 Optical spectroscopy
For each transient, at least one spectrum is usually obtained close to the peak luminosity to establish an initial spectroscopic classification. We primarily utilize the Double Beam Spectrograph (DBSP; Oke & Gunn, 1982) on the Palomar 200-inch Hale telescope and the Spectral Energy Distribution Machine (SEDM; Blagorodnova et al., 2018; Rigault et al., 2019; Kim et al., 2022) on the Palomar 60-inch telescope for this purpose. Other instruments used to obtain classification spectra include: the Alhambra Faint Object Spectrograph and Camera at the Nordic Optical Telescope (NOT; Djupvik & Andersen, 2010), the Spectrograph for the Rapid Acquisition of Transients (SPRAT; Piascik et al., 2014) on the Liverpool Telescope, the Low-Resolution Imaging Spectrometer (LRIS; Oke et al., 1995) on the Keck I telescope. See Das et al. (2025) for details on spectral observations and reductions.
We employ the SuperNova Identification (SNID; Blondin & Tonry, 2007) code for these classifications. For spectra with significant host galaxy contamination, we instead utilize superfit (Howell et al., 2005).
4 Extinction Correction
We corrected for Galactic extinction using the maps provided by Schlafly & Finkbeiner (2011). For the corrections, we use the extinction law described by Cardelli et al. (1989) with a value of .
To estimate the host-galaxy extinction, we use the average color of the sample and measure the host assuming that the SN color exceeding the 1 template is caused by host extinction. See Das et al. (2025) for the intrinsic color template and further details.
5 Analysis
5.1 Observed lightcurve properties
@@
We employed a Gaussian Process (GP) algorithm111https://george.readthedocs.io/ (Ambikasaran et al., 2015) to model the -band and -band lightcurves. The top panel of Figure 1 displays the interpolated -band lightcurve, the middle panel shows its first derivative with respect to time (i.e., the slope, in mag day-1), and the bottom panel shows the time derivative of the slope, denoted as (in mag day-2).
@@
To estimate uncertainties in these observables, we adopt a Monte Carlo approach. We simulate 50 realizations of each lightcurve based on Gaussian noise scaled to the flux errors. For each realization, we recompute the observables using the same criteria. The standard deviation of the resulting distribution of values is used as the uncertainty on the measurements of these observables. The plateau duration is measured as . The optically thick plateau duration (OPTd) is measured as the duration between the explosion epoch and the midpoint of the plateau end and tail onset epochs. To identify ECSN candidates and failed SNe with a sharp drop from the plateau and no nickel mass, we measure the drop in -band magnitude from the plateau end to the first epoch of tail onset when it exists. If a tail does not exist, this is marked as “Limit = yes” in Table A or indicated as a V-shaped triangle in Figure 3. The criteria are summarized in Table 1.
5.2 Measuring photospheric velocities
@@
We derive expansion velocities of the H I line by measuring the minima of the absorption troughs in P-Cygni profiles. A third-order polynomial is fit to the region surrounding the absorption minimum to estimate the velocity. For spectra with low resolution or strong host galaxy contamination, we manually inspect and identify the minimum. Uncertainties are estimated via a Monte Carlo method. We first compute a noise spectrum by subtracting a heavily smoothed version of the original spectrum. The standard deviation of this residual defines the spectral noise. We then generate perturbed realizations by adding Gaussian noise of equal standard deviation to the smoothed spectrum and remeasure the velocity for each realization. The 1 spread in the resulting velocity distribution is adopted as the measurement uncertainty. Only measurements with a standard deviation less than 50% of their velocity are considered reliable and included. The measured velocities and their errors are documented in Table F.
@@
We find that LLIIP SNe exhibit systematically lower H I velocities compared to their higher-luminosity counterparts at all phases. To quantify this trend, we perform a linear fit to the log of the velocity as a function of phase. The best-fit relations are shown in Figure 6. The best-fit for the LLIIP sample is: , and for the full sample: , where is the expansion velocity in km s-1 and is the phase in days since explosion. While both LLIIP SNe and brighter IIP SNe show comparable decline rates, LLIIP SNe exhibit lower velocities. At 50 days post-explosion, the average velocity inferred from these fits is approximately km s-1 for LLIIP SNe and km s-1 for SNe with mag.
5.3 Bolometric Luminosity
@@
The distribution of the plateau bolometric luminosity is shown in the left panel of Figure 4. The full sample has a median peak luminosity of . LLIIP SNe have lower peak luminosities with a median of . We find a strong correlation between the mid-plateau bolometric luminosity () and the nickel mass (Figure 4). The nickel mass is measured by fitting the radioactive tail (see Section 5.4.1). The best-fit relation is
with a Pearson correlation coefficient of and a -value .
5.4 Estimating physical and explosion parameters
In order to estimate progenitor and explosion parameters of the Type IIP sample, we use semi-analytic models Nagy & Vinkó (2016) and radiation hydrodynamical models from Moriya et al. (2023). We also use scaling relations from Goldberg et al. (2019) to constrain the explosion parameters of LLIIP SNe.
5.4.1 Lightcurve fitting to Semi-analytical models
The model in Nagy & Vinkó (2016) is based on a two-component configuration consisting of a uniform dense stellar core and an extended low-mass envelope where the density decreases as an exponential function. Nagy & Vinkó (2016); Jäger et al. (2020) claim that the results from the two-component semi-analytic LC model are consistent with current state-of-the-art calculations for Type II SNe. They are useful to derive estimates or constraints of basic parameters like the explosion energy, ejected mass, and initial radius of the progenitor as well as the amount of synthesized radioactive nickel. These models assume spherical symmetry and do not adequately model the initial transient behavior at early stages (t20 days). Thus, we only fit for lightcurve data 20 days after the explosion. We use the MCMC routine in Jäger et al. (2020) to fit these semi-analytical expressions. The Python code used for these fits will be made available on GitHub after publication. The best-fit lightcurves for the SNe with and without radioactive tails are shown in Figure 7. The nickel mass is measured from the lightcurve tail by fitting for both the nickel mass and the gamma-ray trapping efficiency parameter as free parameters. The priors used for the fits are shown in Table B. We fit the semi-analytic models to Type IIP SNe. After marginalization, we have estimates with confidence intervals (1) for each of these parameters. The distribution of the best-fit parameters and their correlations are shown in Figure 9 and Figure 13. The best-fit values for each SN are listed in Table D. The sample statistics are summarized in Table 3.
5.4.2 Lightcurve fitting to radiation-hydrodynamical Type II model grids

### v2

@@ I Introduction
Low-luminosity Type IIP supernovae (LLIIP SNe), defined here as those with mag, as motivated by prior literature (e.g., Pastorello et al., 2004; Spiro et al., 2014; Reguitti et al., 2021), are particularly intriguing because they are thought to originate from progenitor stars with just enough mass to undergo core collapse. These events likely arise from 8–12 progenitors, as supported by pre-SN progenitor imaging (Maund et al., 2005; Li et al., 2006; Mattila et al., 2008; O’Neill et al., 2019; Van Dyk et al., 2023), light curve simulations (Pumo et al., 2017; Fraser et al., 2011), evolutionary models of low-mass red supergiants (RSGs) (Lisakov et al., 2017), and nebular spectroscopy (Jerkstrand et al., 2018). They provide valuable insight into the lower mass threshold for core collapse, which defines the boundary between stars that form white dwarfs and those that form neutron stars or black holes. LLIIP SNe also serve as key candidates for probing the upper mass limit of super-asymptotic giant branch (sAGB) star formation, exploring whether they can explode as electron-capture supernovae (ECSN) (Nomoto, 1984; Kitaura et al., 2006; Janka et al., 2008; Takahashi et al., 2013; Jones et al., 2013; Hiramatsu et al., 2021b; Wang et al., 2025), and understanding their role in shaping elemental abundances in subsequent generations of stars (Tsujimoto and Shigeyama, 1999; Siess and Pumo, 2006). Furthermore, a typical initial mass function (IMF) predicts that more than 40% of CCSN progenitors fall within the 8–12 range, with 25% within the 8–10 subset (Sukhbold et al., 2016). It has also been suggested that LLIIP SNe may instead originate from more massive RSGs with significant fallback (e.g., Woosley and Weaver, 1995; Zampieri et al., 2003; Moriya et al., 2010). Thus, systematic studies of LLIIP SNe are essential for improving our understanding of the progenitors, explosion properties, and the dust and chemical enrichment contributed by nearly half of all massive stars that undergo core-collapse.
However, to date, no study of the lightcurve properties of LLIIP SNe has been conducted using data from a systematic SN survey. Due to their low luminosity, they are relatively difficult to detect and classify. There are only about a dozen such objects presented in the literature, focused on single-object studies or small heterogeneous samples. Early examples such as SN 1997D, SN 1999eu and SN 1999br (Turatto et al., 1998; Pastorello et al., 2004) revealed the defining characteristics of the class: very low plateau luminosities ( mag), unusually narrow P-Cygni profiles indicating expansion velocities of only 1000–2500 km s-1, extremely small 56Ni masses (– M⊙), and low explosion energies. Subsequent studies of nearby events such as SN 2005cs and SN 2008bk (Pastorello et al., 2009; Mattila et al., 2008; Van Dyk et al., 2012) connected these faint light curves to low-mass red supergiant progenitors (8–12 M⊙) through direct pre-explosion imaging and hydrodynamical modeling. Spiro et al. (2014) modeled a set of underluminous Type IIP SNe and confirmed uniformly low explosion energies; Nakaoka et al. (2018) and Reguitti et al. (2021) found similarly low 56Ni yields and slow expansion velocities for SN 2016bkv and SN 2018hwm, respectively; and studies such as Jäger et al. (2020); Müller-Bravo et al. (2020); Yang et al. (2021); Valerin et al. (2022); Kozyreva et al. (2022); Bostroem et al. (2023); Teja et al. (2024); Dastidar et al. (2025) have recently reported additional LLIIP SN candidates with low luminosities and energetics consistent with either low-mass RSG/sAGB explosions or fallback-suppressed 56Ni production. In contrast to these low-luminosity events, canonical Type II SNe typically reach plateau absolute magnitudes of to mag, have plateau durations of 80–120 days, ejecta masses of 8–18 M⊙, explosion energies of order – erg, and 56Ni yields of 0.01–0.1 M⊙ (e.g., Hamuy, 2003; Anderson et al., 2014; Sanders et al., 2015; Valenti et al., 2016; Martinez et al., 2022b; Fang et al., 2025). These properties are generally consistent with explosions of RSG progenitors with initial masses 12–18 M⊙ (e.g., see Van Dyk, 2025).
@@
In Part I of the series of three papers on LLIIP SNe (Das et al., 2025), we measured the luminosity function and volumetric rate of the largest sample to date of 36 LLIIP SNe (and 330 Type IIP SNe) from the Zwicky Transient Facility Census of the Local Universe (ZTF CLU; Bellm et al., 2019; Graham et al., 2019; Dekany et al., 2020; Masci et al., 2019; De et al., 2020) survey. This paper delves into their lightcurve properties. In this paper, we analyze the observed properties of the multi-band lightcurves of a selected subset of 129 Type II SNe and 16 LLIIP SNe. It is well established that the morphology of Type II SN lightcurves depends on key progenitor properties (e.g., final mass, envelope mass, and radius) and on explosion parameters such as the deposited energy and the amount and distribution of synthesized 56Ni (e.g., Kasen and Woosley, 2009; Goldberg et al., 2019). Here, we measure observed lightcurve parametrs and infer physical and explosion parameters using semi-analytical and radiation-hydrodynamical models that link SN observables to progenitor properties (e.g., Litvinova and Nadezhin, 1983; Hamuy, 2003; Kasen and Woosley, 2009; Goldberg et al., 2019; Morozova et al., 2018; Moriya et al., 2023). Such models have been previously applied to both individual events and larger samples of Type II SNe (Hamuy, 2003; Arcavi et al., 2011; Anderson et al., 2014; Spiro et al., 2014; Sanders et al., 2015; Valenti et al., 2016; Gutiérrez et al., 2017; Morozova et al., 2018; Martinez et al., 2022b, a; Subrayan et al., 2023; Silva-Farfán et al., 2024; Fang et al., 2025). This paper is structured as follows: in Section II, we define the sample selection criteria. The data and extinction correction methods are described in Sections III and IV, respectively. Sections V and VI detail the fitting methods and results. We discuss the implications of the derived physical properties in Section VII and conclude in Section VIII.
II Sample Selection
@@
From the ZTF CLU experiment, we selected Type IIP SNe with well-sampled -band light curves and a clearly defined plateau phase. The selection required (i) a GP-fit peak apparent magnitude of mag, (ii) sufficient temporal coverage around peak and during the plateau (more than ten total detections and at least one detection on both sides of peak, or an equivalent set of conditions ensuring a constrained rise), and (iii) a plateau lasting at least 40 days with mag decline. We refer to Paper I (Das et al., 2025) for the complete sample selection criteria to select 330 Type IIP SNe. To select the sample for lightcurve analysis, we require that the end of the plateau phase is observed in the -band lightcurve. This is to ensure we can extract reliable progenitor and explosion parameters from the lightcurves. The epoch of the end of the plateau () is determined as the epoch where the derivative of the slope of the -band lightcurve () reach a local minimum, and its value is mag day-2, indicating a sharp transition (see Section V.1 and Figure 1 for details). This marks the end of the plateau phase in the lightcurve evolution. Using this criterion, we identified 129 candidates that fulfill this selection criterion. Out of these, 68 objects have -band data points in the radioactive tail phase, while 61 SNe do not have an observed radioactive tail.
III Data
@@
III.1 Optical photometry
We perform forced photometry on the ZTF difference images using the ZTF forced-photometry service developed by Masci et al. (2019) in , and bands. For this work, we consider anything less than a 3 detection an upper limit. In the tail phase, for a few SNe we also took photometry in the , , bands with the Optical Imager (IO:O) at the 2.0 m robotic Liverpool Telescope (Steele et al., 2004) and the Alhambra Faint Object Spectrograph and Camera at the Nordic Optical Telescope (Djupvik and Andersen, 2010). The P60 and LT data sets were processed with the FPipe image subtraction pipeline (Fremling et al., 2016), using Sloan Digital Sky Survey (SDSS; Ahn et al., 2012) and PanSTARRS (PS1; Chambers et al., 2016) reference frames. See Das et al. (2025) for details on photometric observations and reductions. Figure 1 shows a few representative lightcurves as examples. The photometry data and lightcurve plots for all the SNe in the sample will be made available on Zenodo.
III.2 Optical spectroscopy
For each transient, at least one spectrum is usually obtained close to the peak luminosity to establish an initial spectroscopic classification. We primarily utilize the Double Beam Spectrograph (DBSP; Oke and Gunn, 1982) on the Palomar 200-inch Hale telescope and the Spectral Energy Distribution Machine (SEDM; Blagorodnova et al., 2018; Rigault et al., 2019; Kim et al., 2022) on the Palomar 60-inch telescope for this purpose. Other instruments used to obtain classification spectra include: the Alhambra Faint Object Spectrograph and Camera at the Nordic Optical Telescope (NOT; Djupvik and Andersen, 2010), the Spectrograph for the Rapid Acquisition of Transients (SPRAT; Piascik et al., 2014) on the Liverpool Telescope, the Low-Resolution Imaging Spectrometer (LRIS; Oke et al., 1995) on the Keck I telescope. See Das et al. (2025) for details on spectral observations and reductions. We employ the SuperNova Identification (SNID; Blondin and Tonry, 2007) and superfit (Howell et al., 2005) codes for classifications.
IV Extinction Correction
We corrected for Galactic extinction using the maps provided by Schlafly and Finkbeiner (2011). For the corrections, we use the extinction law described by Cardelli et al. (1989) with a value of .
To estimate host-galaxy extinction, we employ the empirical color template constructed in Paper I (Das et al., 2025) from a subset of Type IIP SNe with negligible host reddening (identified via weak or absent host Na I D absorption). The intrinsic template has a 1 scatter of 0.35 mag. For each SN in our sample, we measure the deviation of its observed color from this template and attribute any excess beyond the 1 intrinsic range to host-galaxy reddening, converting it to assuming the same extinction law. The template is available on Zenodo. See Das et al. (2025) for further details.
V Analysis
V.1 Observed lightcurve properties
@@
We employed a Gaussian Process (GP) algorithm111https://george.readthedocs.io/ (Ambikasaran et al., 2015) to fit the -band and -band lightcurves. The top panel of Figure 1 displays the interpolated -band lightcurve, the middle panel shows its first derivative with respect to time (i.e., the slope, in mag day-1), and the bottom panel shows the time derivative of the slope, denoted as (in mag day-2).
@@
To estimate uncertainties in these observables, we adopt a Monte Carlo approach. We simulate 50 realizations of each lightcurve based on Gaussian noise scaled to the flux errors. For each realization, we recompute the observables using the same criteria. The standard deviation of the resulting distribution of values is used as the uncertainty on the measurements of these observables. The plateau duration is measured as . The optically thick plateau duration (OPTd) is measured as the duration between the explosion epoch and the midpoint of the plateau end and tail onset epochs. The distributions of these observed parameters are shown in Figure 2. A sharp luminosity drop between the plateau and the radioactive tail is a key prediction for both ECSNe and failed SNe with very low or negligible 56Ni production (e.g., Tominaga et al., 2013; Turatto et al., 1998). To identify such ECSN and failed-SN candidates in our sample, we measure the drop in the -band magnitude from the end of the plateau to the first epoch on the tail, whenever the tail is detected. If a tail does not exist, this is marked as “Limit = yes” in Table A or indicated as a V-shaped triangle in Figure 3. These photometric definitions follow the same physical interpretation as those adopted by Anderson et al. (2014), who also identify the plateau phase through changes in the light-curve slope and characterize the transition to the radioactive tail using the steep post-plateau decline. Our implementation differs in that we use the derivatives of the lightcurve to define these transitions in a fully automated and objective way, whereas Anderson et al. (2014) relied on parametric spline fits and visually guided measurements for some phases. The data and code used for these measurements will be made publicly available on Zenodo and GitHub upon publication. The criteria are summarized in Table 1.
V.2 Measuring H I velocities
@@
We derive expansion velocities of the H I line by measuring the minima of the absorption troughs in the P-Cygni profiles. A third-order polynomial is fit to the region surrounding the absorption minimum to estimate the velocity. For spectra with low resolution or strong host-galaxy contamination, we manually inspect and identify the minimum. Uncertainties are estimated via a Monte Carlo method. We first compute a noise spectrum by subtracting a heavily smoothed version of the original spectrum; the standard deviation of this residual defines the spectral noise. We then generate perturbed realizations by adding Gaussian noise of equal standard deviation to the smoothed spectrum and remeasure the velocity for each realization. The 1 spread in the resulting velocity distribution is adopted as the measurement uncertainty. Only measurements with a standard deviation less than 50% of their velocity are considered reliable and included. The measured velocities and their errors are documented in Table F.
@@
We find that LLIIP SNe exhibit systematically lower H I velocities compared to their higher-luminosity counterparts at all phases Figure 4. We fit a power-law of the form to the binned velocity measurements. Photospheric velocities in Type II SNe are expected to follow a power-law decline during the plateau phase due to homologous expansion and the recession of the recombination front (e.g., Kasen and Woosley 2009). The best-fit relations are shown in Figure 4. For the LLIIP sample, we obtain , and for the full sample , where is in units of km s-1 and is the phase in days since explosion.
In principle, Fe II is a better tracer of the photospheric velocity than H I (e.g., Dessart and Hillier 2005). However, most of our spectra were obtained with SEDM at , and we do not have sufficient resolution and SNR to robustly measure the Fe II absorption profile in most cases. We therefore use the H velocity as a practical proxy. When needed for comparison with studies that quote Fe II velocities, we employ the empirical scaling relation between H and Fe II derived by Faran et al. (2014),
to estimate the corresponding Fe II velocity. These measured slopes (–) lie within the range of shallow power-law declines inferred for Type II SN photospheric velocities in previous observational studies (e.g., Faran et al., 2014; de Jaeger et al., 2019; Lin et al., 2024). While both LLIIP SNe and brighter IIP SNe show comparable decline rates, LLIIP SNe exhibit lower velocities. At 50 days post-explosion, the average velocity inferred from these fits is approximately km s-1 for LLIIP SNe and km s-1 for SNe with mag. These velocities are broadly consistent with those reported for Type II SNe in the literature (Faran et al., 2014; de Jaeger et al., 2019; Lin et al., 2024).
V.3 Bolometric Luminosity
@@
The distribution of the plateau bolometric luminosity is shown in the left panel of Figure 5. The full sample has a median peak luminosity of . LLIIP SNe have lower peak luminosities with a median of . We find a strong correlation between the mid-plateau bolometric luminosity () and the nickel mass (Figure 5). The nickel mass is measured by fitting the radioactive tail (see Section V.4.1). The best-fit relation is
with a Pearson correlation coefficient of and a -value . These luminosities are consistent with previous measurements for both normal and low-luminosity Type II SNe. Large-sample studies typically find mid-plateau bolometric luminosities of –42.7 erg s-1 for ordinary SNe IIP (e.g., Bersten and Hamuy, 2009; Pejcha and Prieto, 2015; Valenti et al., 2016), whereas underluminous events such as SN 1997D, SN 2005cs, and SN 2008bk exhibit significantly fainter plateaus with –41.7 erg s-1 (Spiro et al., 2014). The tight correlation we find between and is also consistent with earlier work showing that fainter plateau luminosities correspond to smaller 56Ni yields (Hamuy, 2003; Pejcha and Prieto, 2015; Müller et al., 2017; Martinez et al., 2022b).
V.4 Estimating physical and explosion parameters
In order to estimate progenitor and explosion parameters of the Type IIP sample, we use semi-analytic models Nagy and Vinkó (2016) and radiation hydrodynamical models from Moriya et al. (2023). We also use scaling relations from Goldberg et al. (2019) to constrain the explosion parameters of LLIIP SNe.
V.4.1 Lightcurve fitting to Semi-analytical models
The model in Nagy and Vinkó (2016) is based on a two-component configuration consisting of a uniform dense stellar core and an extended low-mass envelope where the density decreases as an exponential function. Nagy and Vinkó (2016); Jäger et al. (2020) claim that the results from the two-component semi-analytic LC model are consistent with current state-of-the-art calculations for Type II SNe. They are useful to derive estimates or constraints of basic parameters like the explosion energy, ejected mass, and initial radius of the progenitor as well as the amount of synthesized radioactive nickel. The model assumes spherical symmetry and do not adequately model the initial transient behavior at early stages (t 20 days). Thus, we only fit for lightcurve data 20 days after the explosion. We use the MCMC routine in Jäger et al. (2020) to fit these semi-analytical expressions. The Python code used for these fits will be made available on GitHub after publication. The nickel mass is measured from the lightcurve tail by fitting for both the nickel mass and the gamma-ray trapping efficiency parameter as free parameters. The priors used for the fits are shown in Table B. We fit the semi-analytic models to Type IIP SNe. After marginalization, we have estimates with confidence intervals (1) for each of these parameters. The best-fit values for each SN are listed in Table D. The sample statistics are summarized in Table 3.
V.4.2 Lightcurve fitting to radiation-hydrodynamical Type II model grids